OpenStudy (anonymous):

Is the expression x3•x3•x3 equivalent to x3•3•3? Why or why not? Explain your reasoning.

OpenStudy (trojanpoem):

Is this x^3 or 3x

OpenStudy (trojanpoem):

if it's x^3 * x^3 * x^3 then = x^9 the other is x^3*3*3 = x^27 not equivalent if it's x * 3 *x * x *3 * x *3 = 27x^3 the other is 27x in both cases not equivalent

OpenStudy (trojanpoem):

Reasoning for the first is that multiplied entities with the same base the power is summed not multiplied

OpenStudy (anonymous):

x'3